OpenStudy (anonymous):

simplify this problem 2(16-13)/3^2+2^2

OpenStudy (blondisaurus):

Distribute the 2 on the top part of the equation

OpenStudy (anonymous):

so it will be 32+26 at the top

OpenStudy (blondisaurus):

Remember that it's minus 6 not plus but yes

OpenStudy (blondisaurus):

minus 26**

OpenStudy (anonymous):

that is 6/9 + 4

OpenStudy (anonymous):

so its 32-26

OpenStudy (blondisaurus):

Yeah, which equals?

OpenStudy (anonymous):

13

OpenStudy (blondisaurus):

32-26 equals 13? Are you sure about that

OpenStudy (anonymous):

is the answer going to be6/13

OpenStudy (blondisaurus):

Lol, im trying to help you solve it

OpenStudy (anonymous):

lol so I was right the first time do I set 32-26=13 then substitute it by adding 26 on both sides

OpenStudy (blondisaurus):

32-26= 6, I dont see where you keep getting 13

OpenStudy (anonymous):

I had squared 3^2+2^ I had added them and got 13 that's wea I got that from

OpenStudy (blondisaurus):

ok so yes the denominator is 13, the numerator is 6

OpenStudy (anonymous):

I got 32/32 so will the answer be 1 becuz any number divided by it self is 1

OpenStudy (blondisaurus):

How did you get 32 on the top?!

OpenStudy (anonymous):

I did 32-26=6 I added 26 on both sides and I brought my 32 down and I added 6+26 I got 32 for thatso 32/32=1

OpenStudy (blondisaurus):

You dont add 26 to anything now bring anything down

OpenStudy (blondisaurus):

32-26=6, that's it. Nothing else can be done

OpenStudy (anonymous):

so that's the answer

OpenStudy (blondisaurus):

The bottom portion is 3^2 thats 9,2^2 equals 4. Add them and thats the bottom portion

OpenStudy (anonymous):

@grudge101

OpenStudy (anonymous):

I got 13

OpenStudy (blondisaurus):

Ok, so the top is 6, the bottom 13. There you go

OpenStudy (anonymous):

that's the answer 6/13

OpenStudy (anonymous):

2(16-13)/3^2 + 2^2 = 2(3)/9 + 4 = 6/9 + 4 = 42/9 = 14/3